how do i use the product property to find log 24
Answer by htmentor(1343)

Using the product rule for logarithms, i.e, log(a*b) = log(a) + log(b), we can write
log(24) = log(4*6) = log(2*2*2*3) = log(2^3*3) = 3log(2) + log(3)